OpenStudy (anonymous):

Multiply. State any restrictions on the variable. 12-x-x^2/x^2-8x+15 Can you guys show me the steps to this? I don't get it

OpenStudy (whpalmer4):

\[\frac{12-x-x^2}{x^2-8x+15}\]Factor the numerator and denominator. Before doing any cancellation, find the values of \(x\) that make the denominator = 0. Those values are the restricted values — \(x\) cannot be any of them as you cannot meaningfully divide by 0.

OpenStudy (whpalmer4):

You have to do this check before you cancel common terms from numerator and denominator because they are still restrictions on the simplified fraction.

OpenStudy (anonymous):

I know it's (x,-3) and (x,-5). What do I do next?

OpenStudy (whpalmer4):

Show me your factored denominator...

OpenStudy (anonymous):

What is the factored denominator?

OpenStudy (whpalmer4):

I asked you to factor the denominator of that fraction. What does it look like after you have factored it?

OpenStudy (anonymous):

I didn't factor it. I don't know how.

OpenStudy (whpalmer4):

You don't know how to factor \[x^2-8x+15\]?

OpenStudy (anonymous):

:l Sorry, I'm stupid.

OpenStudy (whpalmer4):

No, not stupid, but maybe you haven't learned it for some reason. Has it been covered in class? You really can't do this problem without knowing how to factor...

OpenStudy (anonymous):

I dunno

OpenStudy (anonymous):

can you show me the steps? c:

OpenStudy (anonymous):

please?

OpenStudy (whpalmer4):

yes, sorry, had to take care of something

OpenStudy (whpalmer4):

Okay, we have a trinomial (fancy word for 3 terms): \[x^2-8x+15\] We want to split it into the product of two binomials (fancy word for 2 terms): \[x^2-8x+15 = (x+a)(x+b)\]But how?

OpenStudy (anonymous):

it's okay:p

OpenStudy (whpalmer4):

Let's multiply \((x+a)(x+b)\): \[(x+a)(x+b) = x(x+b) + a(x+b) = x*x + x*b + a*x + a*b\]\[=x^2 + (a+b)x + a*b\]Agreed so far?

OpenStudy (whpalmer4):

hello...

OpenStudy (anonymous):

sorrry. This thing wasn't letting me type.

OpenStudy (whpalmer4):

no problem...okay with what I've done so far?

OpenStudy (anonymous):

yes. And if I don't respond, just keep showing me the steps. Cause this thing is getting stupid.

OpenStudy (anonymous):

the typing

OpenStudy (whpalmer4):

so let's compare what we've got: \[x^2+(a+b)x + a*b\]\[x^2-8x+15\]To make those equivalent, we need to pick \(a,b\) such that \[(a+b) = -8\]and\[a*b=15\]Your mission is to find a pair of numbers that make that work.

OpenStudy (whpalmer4):

Hopefully it is obvious that they both have to be negative in this case...

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

how can we get a*b = 15, with both a and b negative? -1*-15 = 15 -3*-5 = 15 -5*-3 = 15 -15*-1 = 15 We also need them to add up to -8 That leaves a=-3, b=-5 (or vice versa) So our factored version is \[(x-3)(x-5)\]Let's check by multiplying that out: \[(x-3)(x-5) = x*x -5*x -3*x-3*(-5) = x^2-5x-3x+15\]\[(x-3)(x-5)=x^2-8x+15\] That means our fraction can be written\[\frac{12-x-x^2}{(x-3)(x-5)}\] Now, to find the restrictions on \(x\), we set the denominator = 0 and solve for the values of \(x\) that make it = 0. We have a product as the denominator, so that makes this task very easy: \[(x-3)(x-5) = 0\]\[x-3=0\]\[x-5=0\]\[x=3,\,x=5\] So our restrictions are that \(x \ne 3, x\ne 5\)

OpenStudy (anonymous):

is that it?

OpenStudy (whpalmer4):

We should factor the numerator (even though we don't need to for this problem) for practice. \[-x^2-x+12\](I reversed the order of the terms, but of course the value does not change) We can factor out a common \(-1\) from each term:\[-1(x^2) -1(x) -1(-12) = -1(x^2+x-12)\]We'll set that \(-1\) aside and just factor the quantity in the parentheses. We'll put the \(-1\) factor back when we are done. \[x^2+x-12 = x^2+1x-12\]Again, we want to find two numbers \(a,b\) such that \(a+b = 1\) and \(a*b = -12\) Let's factor -12: -1*12 -2*6 -3*4 -4*3 -6*2 -12*1 which of those pairs adds up to +1? That will be our factoring. -3+4 = +1 \[x^2+x-12 = (x-3)(x+4)\]Checking: \[(x-3)(x+4) = x*x + 4*x -3*x -3(4) = x^2 + x - 12\] We had a \(-1\) that we factored out, so our fully factored numerator would be: \[-x^2-x+12 = -1(x-3)(x+4)\] Giving us a fully factored fraction of \[\frac{-1(x-3)(x+4)}{(x-3)(x-5)}\] We could then simplify that: \[\frac{-1(x-3)(x+4)}{(x-3)(x-5)}= \frac{-1\cancel{(x-3)}(x+4)}{\cancel{(x-3)}(x-5)} = -\frac{(x+4)}{(x-5)}\] But our restrictions remain the same, even though we no longer have the \((x-3)\) term in the denominator!

OpenStudy (whpalmer4):

That final fraction could also be written as \[-\frac{x+4}{x-5} = \frac{x+4}{5-x}\]

OpenStudy (whpalmer4):

or \[\frac{4-x}{x-5}\]

OpenStudy (whpalmer4):

Now, I only covered factoring trinomials where the coefficient of the \(x^2\) term is 1. Unfortunately, it doesn't have to be 1, and it's more complicated to factor when it isn't. The procedure is sort of similar, and I'm happy to show it to you right now, or give you a rain check and explain it at some other time. Which do you want?

OpenStudy (anonymous):

well do I have everything I need already?

OpenStudy (whpalmer4):

you have the answer to this problem, and I've given you a brief tutorial on how to factor trinomials where the leading coefficient is 1. That may be all you need for now.

OpenStudy (whpalmer4):

By the way, the first word of your problem statement is "multiply", but then you showed a division...we only have the restriction on variable business if it is a division, did you make a mistake when you typed "multiply"?

OpenStudy (anonymous):

no. :/

OpenStudy (whpalmer4):

And the fraction as I typed it, that's what appears in your problem?

OpenStudy (anonymous):

i copied what it said in the book

OpenStudy (whpalmer4):

That's not answering my question...

OpenStudy (whpalmer4):

Does the fraction as I formatted it look like the problem in the book?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

Okay, well, must be an error in the book. Wouldn't be the first time that has happened, or likely the last. Any questions or things that you'd like cleared up?

OpenStudy (anonymous):

nope. thanks do

OpenStudy (whpalmer4):

Did you block me for some reason?